### v2

@@ 5 Gosset’s polytope $4_{21}$
Gosset’s polytope $4_{21}$ is a uniform polytope in eight dimensions with 240 vertices lying on the surface of a 7-sphere. It has 120 rays that form 2025 bases, with each ray occurring in 135 bases, so its ray-basis symbol is $120_{135}-2025_{8}$.
The triacontagonal projection arranges the 120 rays, in sets of fifteen, at the vertices of eight concentric regular pentadecagons which we label by the letters A through H in order of decreasing circumradius; see Table 10 for the numbering of the rays and their coordinates in the plane of the projection.
Pentadecagon Circumradius Angle Rays
A 1.0000 0 1-15
B 0.8135 6 16-30
C 0.6723 6 31-45
D 0.6180 0 46-60
E 0.5027 6 61-75
F 0.4159 6 76-90
G 0.3383 0 91-105
H 0.2091 0 106-120
Table 10: Pentadecagons of Gosset’s polytope, labeled A through H in order of decreasing circumradius. The rays on the pentadecagons are numbered in counter-clockwise order, beginning from the ray lying on or just above the positive x-axis. The starting angles of rays on all pentadecagons are indicated in the third column. The rays on all pentadecagons are spaced $24{{}^{\circ}}$ apart.
The 2025 bases of Gosset’s polytope break up into 135 orbits of fifteen bases each, with the orbits having 33 different basis profiles. Table 2 of the Supplementary Materials lists a generator for each orbit, along with its basis profile. Since there are 135 generators but only 33 profiles, many profiles have more than one generator associated with them.
An examination of the 33 basis profiles shows that the sixteen generators associated with eight of them directly give parity proofs (two of which are shown as the first two entries in Table 12). However, a large number of other parity proofs can be constructed by putting the generators together in the right combinations. We indicate how this can be done using a technique identical to that we used for the 120-cell.
Using the information in Table 2 of the Supplementary Materials, one can set up a 8 x 135 dimensional matrix $M$ whose $i,j$-th element is the number of times pentadecagon $i$ occurs in the basis profile of generator $j$. We used Maple to determine that the nullspace of $M$ has dimension 131 and found a set of linearly independent vectors in it. These vectors, which can be written as words of one to five letters (representing generators), are given in Table 3 of the Supplementary Materials. By combining these words together in all possible ways using the procedure laid out earlier for the 120-cell, and retaining only the odd length words, one gets a total of $2^{130}$ parity proofs for Gosset’s polytope.
We now exhibit some of these proofs. To spare the reader from the necessity of looking up the Supplementary Materials file, we list a few of the generators of Gosset’s polytope in Table 11 and some of the proofs that can be constructed out of them in Table 12. A calculation similar to that in Eq.(2) can be carried out for each of the parity proofs in Table 12 to obtain its ray-basis symbol, if one takes the basis profiles of the relevant generators from Table 11.
@@
Table 11: Shown above are 12 of the 135 generators of Gosset’s polytope, with their letter labels and basis profiles. Each letter of a basis profile indicates the pentadecagon to which the corresponding ray of the generator belongs.
@@
Table 12: Some parity proofs of Gosset’s polytope. The transition from the parity proofs in the first column to the pentadecagon counts in the second and the ray-basis symbols in the third can be made in the way illustrated in Eq.(2) if one replaces each letter of the parity proof by its basis profile as given in Table 11.
@@ A A D D G G H H A A D D G G H H
Table 13: Parity proofs contained in the word $e_{1}e_{2}$. The generators $e_{1}$ and $e_{2}$ are shown at the top of the blocks and the bases they give rise to are shown below them. The bases in the left and right blocks each give a $30_{2}-15_{4}$ parity proof. However, three smaller parity proofs can be obtained by combining bases from both the blocks. One of these proofs is shown in bold font and another in italic font, and they both have the ray-basis symbol $36_{2}-9_{8}$. The third proof, which also has this symbol, has some bases in common with the other two proofs but also has three bases that are its own. If the bases are numbered 1 to 30 going up to down and left to right, the bases associated with the third proof are 1,4,6,9,11,14,17,22 and 27.
We next determine which proofs of Gosset’s polytope are minimal. Since $n=135$ and $k=131$ in this case, it follows from Proposition 1 that only proofs of length $l\leq n-k+1=5$ can be minimal. One finds from the MacWilliams identities that the number of proofs of length 1,3 and 5 are 16, 25812 and a little over 21 million, respectively, with the number growing rapidly after that and going to zero beyond a length of 133 (see Table 4 of the Supplementary Materials for full details). While Proposition 1 asserts categorically that proofs of length 7 or longer are not minimal, it is ambivalent about proofs of length 5 or less. Only a direct examination of any proof of length 5 or less can reveal if it is minimal. We have determined that all the proofs in Table 12 are minimal and irreducible (i.e., they do not have smaller parity proofs in them).
We end by giving an example of a word of two letters that is not a parity proof but still very interesting. The word is $e_{1}e_{2}$, and its bases are listed in Table 13. Each of the letters gives a parity proof of fifteen bases that is both minimal and irreducible. However, the word as a whole is reducible because it has three smaller parity proofs embedded in it, each of which inherits its bases from both the letters. Two of these proofs are shown in bold and italic font, and they have no bases in common. The third proof has some bases in common with the other two proofs in addition to some bases of its own. All three proofs involve nine bases and are the smallest parity proofs in Gosset’s polytope (in that there are no proofs with a smaller number of bases). For comparison, the smallest parity proofs in the 600-cell and 120-cell consist of thirteen and fifteen bases, respectively.
The list of proofs in Table 12 can be extended to proofs of larger length by making use of the data given in Tables 2 and 3 of the Supplementary Materials. Repeated application of Eqs.(1) and (2) to this data allows progressively longer proofs to be constructed.
@@ 6 Discussion
This paper shows how the triacontagonal projections of the 600-cell, 120-cell and $4_{21}$ introduced by Coxeter can be used to obtain parity proofs of the BKS theorem. The special feature of these proofs is that they involve both rays and bases that are multiples of fifteen, as a consequence of their being embedded in a Kochen-Specker diagram derived from a triacontagonal projection. Each proof can be represented by a word made up of distinct letters, each of which represents a set of fifteen bases whose members are related to each other by rotations about the center of the projection. A word makes it possible to write down the ray-basis symbol of the associated proof without first recovering its bases. Many words can be broken down into smaller parity proofs involving fewer bases.
The approach of this paper can be contrasted to other approaches Waegell2011b -Megill that have been used to obtain parity proofs in these polytopes. Parity proofs in all three polytopes were first obtained in Waegell2011b -Waegell2015 by using an algorithm that systematically picked out the proofs, beginning from the smallest and proceeding to larger ones. These methods were improved and generalized in Pavicic2011 and Megill , which made it possible to obtain a much fuller picture of the quantum contextual sets in these polytopes (of which the parity proofs are only a small subset).
A rather different approach was taken in lisonek1 , in which an algebraic method was used to obtain parity proofs in all three polytopes. We used just this method in Sec.4 to obtain the fifteen-fold parity proofs of the 120-cell, but with the difference that whereas lisonek1 considered the incidence matrix between the individual rays and bases, we considered the same matrix between pentadecagons (which are sets of fifteen rays) and generators (which give rise to orbits of fifteen bases). Thus our approach can be considered a coarse-grained version of their fine-grained one. The fine-grained approach yields all the parity proofs in these polytopes, whereas our coarse-grained approach, which applies only to polytopes with a triacontagonal projection, picks out only proofs with a fifteen-fold symmetry about the center of the projection. The proofs obtained with our approach are just a small subset of the proofs obtained using the fine-grained approach. However, as a compensation, we can give a far more detailed account of the proofs in this limited class, as we have tried to illustrate in the results given in the last three sections.
Perhaps the most important message of this paper is that Coxeter’s triacontagonal projections lead directly to parity proofs in all three polytopes if only one realizes that the thirty-fold symmetry of the projection translates into a fifteen-fold symmetry of both the rays and the bases about the center of the projection (which is also the center of the Kochen-Specker diagram). This leads directly to parity proofs of fifteen bases in all three polytopes, without our having to hunt through the bases to find them.
@@
We end by mentioning a couple of open questions about the polytopes studied in this paper. The first is that the largest noncontextual sets in these polytopes are not known. Another way of stating the question is this: What is the largest number of bases with the property that every ray in them can be assigned the value 0 or 1 in such a way that every basis has exactly one ray assigned the value 1 (and all the others assigned the value 0) in it? This question is interesting because the gap between the number of bases in this set and the total number of bases of the polytope provides a measure of how contextual (or nonclassical) the bases of the polytope are. It was shown in BPQS that there is a two-way connection between a quantum contextual set and a perfect quantum strategy in a nonlocal two player game based on that set. Determining the largest noncontextual set would allow us to determine how large a quantum advantage the polytope provides in such a game. Unfortunately, we know of no good (i.e., efficient) algorithm for solving this problem.
@@
A second question pertains to rigidity. The rays (or the equivalent projectors) of a polytope are said to be rigid if any set of projectors (not necessarily of rank one and possibly in a dimension greater than that of the polytope) that satisfy the same orthogonality relations as them can be mapped into them in a one-to-one fashion by a suitable unitary transformation. Proving the rigidity of a set of rays is not an easy task and not many examples of such sets are known. Rigidity is of interest because only a KS set that is rigid can be Bell self-testedMayers -Xu . Therefore it is of interest to add to the known store of rigid KS sets. An important advance in this direction was recently made in Trandafir , where two rigid KS sets were identified in $\mathbb{C}^{3}$, with one of them (of 31 vectors) being conjectured to be the smallest such set in this dimension.
@@
We would like to raise the question of whether the rays of the three polytopes studied here are rigid. Let us first consider the 600-cell. If each ray is considered to be an element of $\mathbb{RP}^{3}$, it is determined by three real numbers and the 60 rays are determined by 180 real numbers. However the 75 bases of the 600-cell imply 450 orthogonality relations between its rays, and since the number of constraints imposed by orthogonality (450) greatly exceeds the number of ray components (180), it would seem that the rays are uniquely determined. However, there could be rays in $\mathbb{RP}^{n}$ or $\mathbb{CP}^{n}$ for $n>3$ that obey the same orthogonality relations as the rays of the 600-cell but are not unitarily equivalent to them, which would demonstrate that the 600-cell is not rigid. We show in the Appendix that of a subset of the rays of $E_{8}$ (or Gosset’s polytope) satisfy all the orthogonality relations of the rays of the 600-cell but are not unitarily equivalent to them, thereby establishing that the 600-cell is not rigid.
@@
However we cannot make a statement about the rigidity of either the 120-cell or Gosset’s polytope. The number of orthogonalities in these polytopes (particularly the latter) is so large and imposes so many constraints that we feel that rigidity may well be forced. However, this is just a conjecture and we have not come close to proving it. We leave it as an open problem to be considered. If these polytopes are rigid, that would certainly enhance their interest as far as quantum contextuality is concerned.
@@
One final point should be made. The bases of all three polytopes can be obtained as the products of powers of three unitary operators acting on the computational basisWaegell2011b -Waegell2015 . If gates corresponding to the basic unitary operators can be designed, it would be possible to realize these polytopes (i.e., their bases) experimentally in the laboratory. This might be worth keeping in mind for the time when the technology for doing it becomes available.
@@
We follow Moody and Patera Moody in obtaining the 240 root vectors of $E_{8}$ (or the vertices of Gosset’s polytope) using icosians, which are quantities related to the symmetries of the regular icosahedron. We first construct the 120 root vectors of $H_{4}$ (which comprise the vertices of a regular 600-cell) in two different ways. Let $\beta=(1+\sqrt{5})/2$ be the golden ratio and $\alpha=(1-\sqrt{5})$ its conjugate. Next, let $\mathbb{Z}[\alpha]=\{m+n\alpha:m,n\in\mathbb{Z}\}$ denote the “golden ring” consisting of all linear combinations of 1 and $\alpha$ with integer coefficients. In particular, we notice there is a bijection $\phi:\mathbb{Z}[\alpha]\rightarrow\mathbb{Z}^{2}$ which maps $m+n\alpha$ to $(m,n)$ and that this bijection extends to a map $\phi:\mathbb{Z}[\alpha]^{4}\rightarrow\mathbb{Z}^{8}$.
Let $G\subset GL(4,\mathbb{R})$ be the linear group generated by all even permutations and arbitrary sign changes of the four coordinates of a vector $(w,x,y,z)\in\mathbb{R}^{4}$. This group has $|G|=12\cdot 16=192$ elements. The roots of $H_{4}$ can be constructed as the union of the orbits of $(2,0,0,0)$, $(1,1,1,1)$ and $(0,\alpha,1,\beta)=(0,\alpha,1,-\alpha+1)$ under the action of $G$. The 120 vectors thus obtained represent the vertices of a 600-cell on a sphere of radius 2. The vectors come in pairs that are the negatives of each other, and if we keep just one member of each pair we get the 60 rays of the 600-cell. Let us call this set of 60 vectors (or rays) $H_{4}^{a}$. If we multiply each of the vectors of $H_{4}^{a}$ by $\alpha$ we get a set $H_{4}^{b}$ which serves as the set of vertices of a second 600-cell, scaled by a factor of $\alpha$ relative to the first. Analogous to $H_{4}^{a}$, we may represent the respective orbits by $(2\alpha,0,0,0)$, $(\alpha,\alpha,\alpha,\alpha)$ and $(0,1+\alpha,\alpha,1)$.
@@
Acknowledgement. GNP was supported by the Japanese Society for the Promotion of Science (JSPS) as a JSPS post-doctoral fellow and by the JSPS KAKENHI Grant number 24KF0176.

### v3

@@ 6 Gosset’s polytope $4_{21}$
Gosset’s polytope $4_{21}$ is a uniform polytope in eight dimensions with 240 vertices lying on the surface of a 7-sphere. Its vertices come in antipodal pairs, so it has 120 rays. The rays form 2025 bases, with each ray occurring in 135 bases, so the ray-basis symbol of this polytope is $120_{135}-2025_{8}$.
The Kochen-Specker diagram obtained from the triacontagonal projection arranges the rays at the vertices of eight concentric regular pentadecagons which we label by the letters A through H in order of decreasing circumradius. See Table 11 for the numbering scheme we use for the rays.
Pentadecagon Rays
A 1-15
B 16-30
C 31-45
D 46-60
E 61-75
F 76-90
G 91-105
H 106-120
Table 11: Pentadecagons of Gosset’s polytope, labeled A through H in order of decreasing circumradius. The rays are numbered in counterclockwise order around the pentadecagons, beginning from the ray lying on or just above the positive x-axis.
The 2025 bases of Gosset’s polytope break up into 135 orbits of fifteen bases each [^10] If one numbers the rays as in Table 11 and uses triacontagonal coordinates for them, it becomes easy to pick out the generators of the bases and also to visualize the bases in the projection plane., with the orbits having 33 different basis profiles. Table 2 of the Supplementary Materials lists a generator for each orbit, along with its basis profile. Since there are 135 generators but only 33 basis profiles, many of the profiles have more than one generator associated with them.
An examination of all the basis profiles shows that sixteen generators associated with eight of the profiles directly give rise to parity proofs (two of which are shown as the first two entries in Table 13). However, many other proofs can be constructed by putting the generators together in the right combinations. This can be done by a method similar to that we used for the 120-cell in Sec.4, but with the matrix $M$ now having dimensions 8 x 135. The matrix can be set up by using the information about the generators and their basis profiles given in Table 2 of the Supplementary Materials. A calculation shows that the nullspace of $M$ mod 2 has dimension 131 and a set of linearly independent vectors in it are shown as the words of one to five letters (representing generators) in Table 3 of the Supplementary Materials. By combining these words together in all possible ways using the rule given in Eq.(2) and keeping only the odd length words, one can generate all the $2^{130}$ fifteen-fold parity proofs in this polytope.
We now exhibit a few of the proofs. To spare the reader the necessity of looking up the Supplementary Materials, we have listed a few of the generators in Table 12 and some of the proofs that can be constructed from them in Table 13.
@@
Table 12: Generators of Gosset’s polytope. The first column shows 12 of the 135 generators, along with their letter labels, with their basis profiles shown to their right. The subscripts on the letter labels are used to distinguish generators belonging to the same basis profile.
@@
Table 13: Fifteen-fold proofs of Gosset’s polytope. The transition from the words/proofs in the first column to the ray-basis symbols in the third can be made as in Eqs.(1) and (3) by using the basis profiles given in Table 12. All the proofs shown here are irreducible (i.e., they do not have any smaller parity proofs in them.
@@ A A D D G G H H A A D D G G H H
Table 14: Parity proofs in the word $e_{1}e_{2}$. The generators $e_{1}$ and $e_{2}$ are shown at the top of the left and right columns and their orbits of bases are shown below them. The bases in the left and right blocks each give a $30_{2}-15_{4}$ parity proof. However, there are three smaller parity proofs that get their bases from both columns of the table. One of them is shown in bold font and another in italic font. These two proofs have no bases in common and they both have the ray-basis symbol $36_{2}-9_{8}$. The third proof, which also has the same symbol, has some bases in common with the other two proofs but also three bases that are its own. If the bases are numbered 1 to 30 from up to down and left to right, the bases associated with the third proof are 1,4,6,9,11,14,17,22 and 27.
Table 14 shows all the bases associated with the two letter word $e_{1}e_{2}$, which is not a parity proof but very interesting all the same. The bases associated with the letters $e_{1}$ and $e_{2}$ are shown in the two halves of the table, and the fifteen bases of each give a parity proof. However there are three smaller parity proofs in the table, each of which inherits its bases both from $e_{1}$ and $e_{2}$ (see table caption for details). All three proofs consist of just nine bases each and are the smallest parity proofs in Gosset’s polytope. This example, like some of those in Table 9, illustrates the fact that words of both even and odd lengths can sometimes have much smaller parity proofs (of a less than fifteen-fold symmetry) embedded in them.
To round out the very limited view we have given of fifteen-fold proofs in Gosset’s polytope, we used the MacWilliams identities to calculate the number of such proofs of various word lengths and found that the number of proofs of word length 1,3 and 5 is 16, 25812 and a little over 21 million, respectively, with the number growing rapidly after that and then falling off and going to zero beyond a word length of 133 (see Table 4 of the Supplementary Materials for more details). The proofs have a great variety of taxonomies, as reflected in their ray-basis symbols.
We finally address the question of which of all the fifteen-fold proofs might be reducible. We can answer this question by using the fact that the fifteen-fold proofs of Gosset’s polytope are like the codewords of an $[n,k,d]$ linear binary code with $n=135$ and $k=131$. Then it follows from Proposition 1 that all proofs with a word length of 7 or larger are reducible and that only the proofs with a word length of five or less might be irreducible. As we pointed out above, there are a little over 21 million proofs of five letters and we know of no way, short of an individual examination of each case, of telling how many of these might be irreducible.
@@ 7 Discussion
We have shown how the triacontagonal projections of the 600-cell, 120-cell and Gosset’s polytope can be used to obtain parity proofs of the BKS theorem having a fifteen-fold symmetry about the center of a suitably chosen Kochen-Specker diagram. The smallest such proofs in all three polytopes, which consist of fifteen bases, follow immediately. However there are many larger proofs that all involve both rays and bases that are multiples of fifteen. We have shown that all such proofs can be written as words made up of an odd number of distinct letters from which all their properties, including the bases of which they are made up, can be inferred. Since the proofs number in the billions and many of them consist of hunderds of bases, the benefits of such an approach are obvious. While our construction captures only the proofs of highest symmetry in all three polytopes, it allows them to be categorized and accessed in a way that was not possible before.
The proofs studied in this paper all have a $C_{15}$ symmetry, which they inherit from the $C_{30}$ symmetry of the triacontagonal projection. The automorphism groups of all the polytopes are actually much larger [^11] The symmetry groups of the 600-cell and 120-cell are identical, while that of Gosset’s polytope is different. The symmetry groups of all three polytopes are discussed in Cox1 ., but none of them has any rotational symmetries larger than $C_{30}$. However, since $C_{5}$, $C_{3}$ and trivial symmetry are all subgroups of $C_{15}$, these polytopes have many parity proofs with these smaller symmetries as well.
Aside from the three polytopes studied in this paper, we are aware of a family of eight-dimensional polytopes, all closely related to Gosset’s polytope, for which triacontagonal projections exist. These polytopes can all be described by the Coxeter diagram below
by replacing one or more of its nodes by ringed nodes in all possible ways. Since this can be done in $2^{8}-1=255$ ways, there are 255 polytopes that can be obtained in this way, of which Gosset’s polytope is one. However the remaining 254 polytopes also possess triacontagonal projections and any bases formed by their rays will therefore occur in orbits of fifteen. Thus it is possible that they also harbor parity proofs similar to those we have found in Gosset’s polytope. However, working out the triacontagonal coordinates of these polytopes and seeing if this expectation is satisfied requires a lot of computation, and we leave it as a problem for the future.
The fifteen-fold symmetry of the triacontagonal projection that we exploited in this paper suggests that a similar approach could prove useful in other cases. We mention two problems that could be tackled with this approach. First, consider the system of 60 rays and 105 bases arising from the triads of commuting observables of the two-qubit Pauli group. This system has been shown to contain a large number of parity proofs Waegell2011 ; lisonek1 , but only a few of them have been exhibited and no trends underlying them have been detected. It is possible that by looking at the automorphism group of this system and its subgroups, one could classify the proofs into different symmetry classes and achieve a much better understanding of them, similar to what has been done here. A far more challenging problem would be to apply this approach to the binary and ternary Golay codes, both of which are known to contain contextuality proofs that are a variant of parity proofsWaegell2022 . However, only one such proof has been identified in each case and neither is particularly small. The large dimensionalities of these systems (24 for the binary code and 12 for the ternary one) and the huge number of bases involved make the usual types of search methods impractical. However it is possible that by looking at the symmetry groups of these systems and identifying the right subgroups, one might be able to uncover at least a subclass of the contextuality proofs in each of them. Just as the triacontagonal projection yielded the smallest fifteen-fold proofs in all three polytopes trivially, it is our hope that symmetry could lead to a similar benefit in these far more intractable cases.
@@
We end by mentioning two open questions about the polytopes studied in this paper. The first concerns the size of the largest noncontextual sets in them, i.e., the largest number of bases with the property that every ray in them can be assigned a 0 or a 1 in such a way that every basis has exactly one ray assigned the value 1 in it. This question is interesting because the gap between the number of bases in the largest noncontextual set and the total number of bases of the polytope provides a measure of how contextual (or nonclassical) the bases of the polytope are. It was shown in BPQS that there is a two-way connection between a quantum contextual set and a perfect quantum strategy in a nonlocal two player game based on that set. Determining the largest noncontextual set in a polytope would convey an idea of how large a quantum advantage it would provide in such a game. Unfortunately, we know of no good (i.e., efficient) algorithm for solving this problem and would welcome any progress that could be made in this direction.
@@
The second question concerns rigidity. The rays (or the equivalent projectors) of a polytope are said to be rigid if any set of projectors (not necessarily of rank one and possibly in a dimension greater than that of the polytope) that satisfy the same orthogonality relations as them can be mapped into them in a one-to-one fashion by a suitable unitary transformation. Proving the rigidity of a set of rays is not an easy task and not many examples of such sets are known. Rigidity is of interest because only a KS set that is rigid can be Bell self-testedMayers -Xu . The discovery of two new rigid KS sets in $\mathbb{C}^{3}$ was recently reported in Trandafir , with one of them (consisting of 31 vectors) conjectured to be the smallest in three dimensions.
@@
We would like to raise the question of whether the rays of the three polytopes studied here are rigid or not. First consider the 600-cell. Each ray is an element of $\mathbb{RP}^{3}$ and determined by three real numbers, and all 60 rays by 180 real numbers. However, the 75 bases of the 600-cell imply 450 orthogonality relations between its rays, and since the number of constraints imposed by orthogonality (450) greatly exceeds the number of ray components (180), it would seem that the rays are uniquely determined by orthogonality. However, there could be rays in $\mathbb{RP}^{n}$ or $\mathbb{CP}^{n}$ for $n>3$ that obey the same orthogonality relations as the rays but are not unitarily equivalent to them, and this would demonstrate that the 600-cell is not rigid. We show in the Appendix that a subset of the rays of Gosset’s polytope satisfy all the orthogonality relations of the rays of the 600-cell, but are not unitarily equivalent to them, and this establishes that the 600-cell is not rigid.
@@
However, we cannot make a statement about the rigidity of either the 120-cell or Gosset’s polytope. The number of orthogonalities in both these polytopes (particularly the latter) is so large and imposes so many constraints that we feel that rigidity might well be forced. However, this is just a conjecture on our part and we have not come close to proving it. We leave it as an open problem to be considered. If either of these polytopes could be shown to be rigid, that would certainly enhance its interest as far as quantum contextuality is concerned.
@@
One final point should be made. The bases of all three polytopes can be obtained as the products of powers of three unitary operators acting on the computational basisWaegell2011b -Waegell2015 . If gates corresponding to the basic operators can be constructed, it would be possible to realize these polytopes (i.e., their bases) experimentally in the laboratory. This might be worth keeping in mind for the future time when it becomes possible.
@@
We follow Moody and Patera Moody in obtaining the 240 root vectors of $E_{8}$ (or the vertices of Gosset’s polytope) using icosians, which are quantities related to the symmetries of the regular icosahedron. We first construct the 120 root vectors of $H_{4}$ (which comprise the vertices of a regular 600-cell) in two different ways. Let $\beta=(1+\sqrt{5})/2$ be the golden ratio and $\alpha=(1-\sqrt{5})/2$ its conjugate. Next, let $\mathbb{Z}[\alpha]=\{m+n\alpha:m,n\in\mathbb{Z}\}$ denote the “golden ring” consisting of all linear combinations of 1 and $\alpha$ with integer coefficients. In particular, we notice there is a bijection $\phi:\mathbb{Z}[\alpha]\rightarrow\mathbb{Z}^{2}$ which maps $m+n\alpha$ to $(m,n)$ and that this bijection extends to a map $\phi:\mathbb{Z}[\alpha]^{4}\rightarrow\mathbb{Z}^{8}$.
Let $G\subset GL(4,\mathbb{R})$ be the linear group generated by all even permutations and arbitrary sign changes of the four coordinates of a vector $(w,x,y,z)\in\mathbb{R}^{4}$. This group has $|G|=12\cdot 16=192$ elements. The roots of $H_{4}$ can be constructed as the union of the orbits of $(2,0,0,0)$, $(1,1,1,1)$ and $(0,\alpha,1,\beta)=(0,\alpha,1,-\alpha+1)$ under the action of $G$. The 120 vectors thus obtained represent the vertices of a 600-cell on a sphere of radius 2. The vectors come in pairs that are the negatives of each other, and if we keep just one member of each pair we get the 60 rays of the 600-cell. Let us call this set of 60 vectors (or rays) $H_{4}^{a}$. If we multiply each of the vectors of $H_{4}^{a}$ by $\alpha$ we get a set $H_{4}^{b}$ which serves as the set of vertices of a second 600-cell, scaled by a factor of $\alpha$ relative to the first. Analogous to $H_{4}^{a}$, the vectors of $H_{4}^{b}$ can be obtained as the union of the orbits of $(2\alpha,0,0,0)$, $(\alpha,\alpha,\alpha,\alpha)$ and $(0,1+\alpha,\alpha,1)$.
@@
Acknowledgements. We would like to thank both the anonymous referees for the many constructive suggestions they made for improving the manuscript. GNP was supported by the Japanese Society for the Promotion of Science (JSPS) as a JSPS post-doctoral fellow and by the JSPS KAKENHI Grant number 24KF0176.